OpenStudy (anonymous):

Solve 9x + 5y + 3z = 4 6x + 2y + 4z = 2 -3x + 3y – 7z = 2 using matrices

OpenStudy (anonymous):

Okay, put it into a 3x3 matrix like this: \[\left[\begin{matrix}9 & 5 & 3\\ 6 & 2 & 4\\ -3 & 3 & -7 \end{matrix}\right]\left(\begin{matrix}x \\ y \\ z\end{matrix}\right)\ = \left(\begin{matrix}4 \\ 2 \\ 2\end{matrix}\right)\]

OpenStudy (anonymous):

and then?

OpenStudy (anonymous):

then express it like so, AB=C with A being the 3x3, B being the variable, and C being the RHS. Then Apply A^-1 to both sides, so we know the solution of x, y, z is A^(-1)B or \[B= A^{-1}C\]

OpenStudy (anonymous):

finding A^(-1) is hard to explain, and very lengthy. I will try and find you a good website for it. This one is quite good, and if you give me a minute I will find the inverse of this one for you to check it against.

OpenStudy (anonymous):

http://www.wikihow.com/Inverse-a-3X3-Matrix

OpenStudy (anonymous):

Okay I'll try and solve this. But help me out with the answer though. I appreciate it.